OpenStudy (anonymous):

i don't understand what to do here: Find ds/dt when θ=π if s=cosθ and dθ/dt=-6...

OpenStudy (anonymous):

Message was truncated. Retry the question.

OpenStudy (sgadi):

given \[s=\cos\theta\] so, we can say \[\frac{ds}{dt}=-\sin\theta \frac{d\theta}{dt}\] also given that \[\theta = \pi, \frac{d\theta}{dt}=-6\] so we get \[\frac{ds}{dt}=-\sin\theta \frac{d\theta}{dt}=0\]